Delsing Canning Company is considering an expansion of its facilities. Its current income statement is as follows:

Sales $ 6,600,000
Variable costs (50% of sales) 3,300,000
Fixed costs 1,960,000
Earnings before interest and taxes (EBIT) $ 1,340,000
Interest (10% cost) 520,000
Earnings before taxes (EBT) $ 820,000
Tax (35%) 287,000
Earnings after taxes (EAT) $ 533,000
Shares of common stock 360,000
Earnings per share $ 1.48

The company is currently financed with 50 percent debt and 50 percent equity (common stock, par value of $10). In order to expand the facilities, Mr. Delsing estimates a need for $3.6 million in additional financing. His investment banker has laid out three plans for him to consider:

  1. Sell $3.6 million of debt at 12 percent.
  2. Sell $3.6 million of common stock at $30 per share.
  3. Sell $1.80 million of debt at 11 percent and $1.80 million of common stock at $40 per share.

Variable costs are expected to stay at 50 percent of sales, while fixed expenses will increase to $2,460,000 per year. Delsing is not sure how much this expansion will add to sales, but he estimates that sales will rise by $1 million per year for the next five years

Delsing is interested in a thorough analysis of his expansion plans and methods of financing.He would like you to analyze the following:

a. The break-even point for operating expenses before and after expansion (in sales dollars). (Enter your answers in dollars not in millions, i.e, $1,234,567.)

Break-Even Point Before expansion After expansion

b. The degree of operating leverage before and after expansion. Assume sales of $6.6 million before expansion and $7.6 million after expansion. Use the formula: DOL = (STVC) / (STVC − FC). (Round your answers to 2 decimal places.)

Degree of Operating Leverage Before expansion After expansion

c-1. The degree of financial leverage before expansion. (Round your answer to 2 decimal places.)

Degree of financial leverage

c-2. The degree of financial leverage for all three methods after expansion. Assume sales of $7.6 million for this question. (Round your answers to 2 decimal places.)

Degree of Financial Leverage 100% Debt 100% Equity 50% Debt & 50% Equity

d. Compute EPS under all three methods of financing the expansion at $7.6 million in sales (first year) and $10.5 million in sales (last year). (Round your answers to 2 decimal places.)

Earnings per Share First Year Last Year 100% Debt 100% Equity 50% Debt & 50% Equity
Answer #1

Ans

a) The break-even point for operating expenses before and after expansion (in sales dollars)

At break even before expansion:

PQ = FC + VC

where PQ equals sales volume at break-even point

Sales = Fixed costs + Variable costs

(Variable costs = 50% of sales)

Sales = $1,960,000 + 0.50 Sales

0.50 Sales = $1,960,000

Sales = $3,920,000

At break even after expansion:

PQ = FC + VC

where PQ equals sales volume at break-even point

Sales = Fixed costs + Variable costs

(Variable costs = 50% of sales)

Sales = $2,460,000 + 0.50 Sales

0.50 Sales = $2,460,000

Sales = $4,920,000

b)

Degree of operating leverage, before expansion, at sales of $ 6,600,000

DOL = Q ( P - VC) / Q ( P - VC) - FC = S - TVC / S - TVC - FC

DOL = $6,600,000 - $ 3,300,000 / $6,600,000 - $ 3,300,000 - $ 1,960,000

DOL = $ 3,300,000 / $ 1,340,000 = 2.46

Degree of operating leverage, before expansion, at sales of $ 7,600,000

DOL = Q ( P - VC) / Q ( P - VC) - FC = S - TVC / S - TVC - FC

DOL = $7,600,000 - $ 3,800,000 / $7,600,000 - $ 3,800,000 - $ 2,460,000

DOL = $ 3,800,000 / $ 1,340,000 = 2.84

c-1) The degree of financial leverage before expansion.

DFL = EBIT / (EBIT - Interest)

DFL = $ 1,340,000 / ($ 1,340,000 - $ 520,000)

DFL = $ 1,340,000 / $ 820,000 = 1.63

c-2. The degree of financial leverage for all three methods after expansion

100% Debt (1) 100% Equity (2) 50% Debt + 50% Equity (3)
Sales $7,600,000 $7,600,000 $7,600,000
- TVC 3,800,000 3,800,000 3,800,000
- FC 2,460,000 2,460,000 2,460,000
EBIT $1,340,000 $1,340,000 $1,340,000
Interest (old) 520,000 520,000 520,000
Interest (new) 432,000 0 198,000
Total Interest (I) $952,000 $520,000 $718,000

DFL = EBIT / (EBIT - Interest)

1. $ 1,340,000 / ( $ 1,340,000 - $ 952,000) = 3.45

2. $ 1,340,000 / ( $ 1,340,000 - $ 520,000) = 1.63

3. $ 1,340,000 / ( $ 1,340,000 - $ 718,000) = 2.15

d)Compute EPS under all three methods of financing the expansion at $7.6 million in sales (first year) and $10.5 million in sales (last year).

100% Debt (1) 100% Equity (2) 50% Debt + 50% Equity (3)
EBIT $1,340,000 $1,340,000 $1,340,000
Total Interest (I) 952,000 520,000 718,000
EBT 388,000 820,000 622,000
Taxes (35%) 135,800 287,000 217,700
EAT $252,200 $533,000 $404,300
Shares (old) 360,000 360,000 360,000
Shares (new) 0 120,000 45,000
Total shares 360,000 480,000 405,000
EPS (EAT /Total shares) $0.70 $1.11 $1.00
100% Debt (1) 100% Equity (2) 50% Debt + 50% Equity (3)
Sales $10,500,000 $10,500,000 $10,500,000
- TVC 5,250,000 5,250,000 5,250,000
- FC 2,460,000 2,460,000 2,460,000
EBIT $2,790,000 $2,790,000 $2,790,000
Total Interest (I) 952,000 520,000 718,000
EBT 1,838,000 2,270,000 2,072,000
Taxes (35%) 643,300 794,500 725,200
EAT $1,194,700 $1,475,500 $1,346,800
Shares (old) 360,000 360,000 360,000
Shares (new) 0 120,000 45,000
Total shares 360,000 480,000 405,000
EPS $3.32 $3.07 $3.33